OpenStudy (anonymous):

Prove using trig: cscx cosx tanx = 1

OpenStudy (solomonzelman):

lets put up some rules before doing anything. \(\LARGE\color{blue}{ \bf csc(x) = \frac{1}{sin(x)} }\) \(\LARGE\color{green}{ \bf \frac{cos(x)}{sin(x)}=cot(x) }\) \(\LARGE\color{red}{ \bf cot(x) \times tan(x)=1 }\)

OpenStudy (solomonzelman):

I got disconnected.

OpenStudy (solomonzelman):

\[\csc(x) \cos(x) \tan(x) = 1\] now lets apply the blue rule I posted. \[\color{blue} { \csc(x) } \cos(x) \tan(x) = 1\] \[\color{blue} { \frac{1}{\sin(x)} } \cos(x) \tan(x) = 1\] \[\color{blue} { \frac{1}{\sin(x)} } \color{blue} { \cos(x) } \tan(x) = 1\] \[\color{blue} { \frac{\cos(x)}{\sin(x)} } \tan(x) = 1\] Do you need help applying the green and red rules I posted or not ?

OpenStudy (anonymous):

I would like help with those also yes. Im new to this section because I was absent so I need the answers so I can review it to learn it. Thanks so much

OpenStudy (solomonzelman):

cos(x) / sin(x) turns into cot(x) and cot(x)tan(x) =1